OpenStudy (anonymous):

64 and 48 are the only degrees in a triangle, the who degree for the triangle is 180. What is the thrid degree so it can equal 180? please help! :)

OpenStudy (anonymous):

By triangle law 64 + 48 + x = 180 Therefore x = 180 - 64 -48 = 180 - 112= 68 Therefore 3rd angle = 68

OpenStudy (anonymous):

thank you so much! :)

OpenStudy (anonymous):

Welcome :)